# MIRACLE II: Unveiling the multi-phase gas interplay in the circumnuclear region of NGC 1365 via multi-cloud modeling

### v3

@@ 3.5 Photoionization modeling with HOMERUN
We model the full set of optical to mid-IR emission using state-of-the-art photoionization models. Specifically, we employed the Highly Optimized Multi-cloud Emission-line Ratios Using photo-ionizatioN (HOMERUN) modeling framework developed by Marconi et al. (2024).
3.5.1 Model grids of AGN and SF
@@
Specifically, the shape of the AGN ionizing radiation field is described by a power law with UV slope , an exponential cutoff exp(-h/k TMax) and the X-ray component slope of =-1.0 linked to the UV through the . We have computed models with following combinations of log(TMax/K) = 4.0, 4.5, 5.0, 5.5, 6.0, 6.5, 7.0 and .
As described in Marconi et al. (2024), the incident spectra for the SF model grids are stellar population models from BPASS v2.3 (Byrne et al., 2022; Stanway & Eldridge, 2018) including binary stellar evolution. These models use a Kroupa (2001) initial mass function and an upper mass cut-off of 300 M⊙. We select models with solar value for the stellar metallicity log(Zstar)=-1.7 and ages in log(age/Myr) of 6.0, 6.4, 6.6, 7.1, 8.8.
The other model parameters are identical for the two suites of models (AGN and SF), namely the ionization parameter log(U) (i.e., the number of ionizing photons compared to that of atoms of neutral hydrogen) ranges from to in step of 0.5 and the hydrogen density of the gas, log(nH / cm-3) from 0 to 7 in steps of 1. For the first iteration HOMERUN spans metallicities log(Z/Z⊙) from -1.0 to 0.4 in steps of 0.2. After this first iteration, HOMERUN performs a final refinement in steps 0.02 on predictions interpolated across the finer grid.
3.5.2 HOMERUN modelling approach
@@
As explained in detail in Marconi et al. (2024), the fitting process starts by selecting a grid of CLOUDY models, defined by hydrogen density () and ionization parameter () for a fixed ionizing spectrum (), gas phase oxygen abundance (), and elemental abundance ratios (). For each grid, the code searches for the best-fitting multi-cloud model by combining individual single-cloud templates. This is achieved by solving a Non-Negative Least Squares (NNLS) problem, where the weights of each cloud are constrained to be non-negative. The procedure is conceptually analogous to stellar population fitting methods (e.g., pPXF; Cappellari & Emsellem 2004), where a galaxy spectrum is reproduced as a linear combination of stellar templates with positive weights. No regularization is applied in the minimization, allowing for full flexibility in the weight distribution. The process is repeated across all available grids, spanning different ionizing continua, metallicities, and abundance scalings. The optimal solution is identified as the one minimizing the loss function , which quantifies the deviation between model predictions and observed emission-line fluxes. This loss function is effectively a reduced statistic, so values indicate a statistically good fit.
When exploring different metallicities, elemental abundances were scaled from the solar photospheric values of Asplund et al. (2021), except for carbon and nitrogen, which were rescaled following the prescriptions of Nicholls et al. (2017). During the fit, all elemental abundances except oxygen were allowed to vary from their initial values. Scaling factors were used for the AGN and SF components. Since the AGN component was assumed to be dust-free while the SF component includes dust, refractory elements can have different scaling factors (due to depletion), whereas non-refractory elements share the same scaling factor in both components.
@@ 3.5.3 HOMERUN fitting results
We fit the fluxes of 60 ionized and neutral atomic emission lines measured from the MIRI and MUSE spectra (32 from MIRI and 28 from MUSE, see Figs. 2 and 9) extracted within a 1.5″ radius aperture centered on the outflow region (see Fig. 6). This aperture was chosen as the largest circular region that is fully covered by all MIRI channels, ensuring the detection of the maximum number of emission lines. The unique approach of HOMERUN enables us to reproduce at the same time 60 different emission lines in different ionisation stages, with the majority of them reproduced with better than 15 accuracy. The data are best matched by a SF component with mean = 3.3, / cm-3) = 1.3, and = -3.3 and an AGN component with mean = 1.4, / cm-3) = 2.9, and = -1.4 (see Fig. 7). Being results of ionized gas region, in the following we will consider .
@@
We find that the SF component is more dust-attenuated than the outflow, in agreement with the extinction map obtained from the Balmer decrement (see Fig. 12) where the higher AV values trace the circumnuclear ring and the nucleus. Interestingly, the difference in AV between SF and AGN components is in line with the median values found in MAGNUM galaxies dividing the emission in disk and outflow (Mingozzi et al., 2019). In Section 3.4, we computed the electron density from the mid-IR [Ne V] and the optical [S II] line doublet ratios within the same 1.5″ radius aperture. The density derived from [Ne V], /cm-3) 2.87, is consistent with the average value of the AGN model components, /cm-3) = 2.9. This is expected since HOMERUN attributes the entire emission of [Ne V] to the AGN (see Table 4), as stars may not produce hard-enough radiation to account for the emission of this transition. Conversely, the [S II] lines yield a relatively higher density of log(ne/cm-3) 2.5 within the 1.5″ radius aperture, compared to log(ne/cm-3) 1.3 from HOMERUN. This is because the average density of the SF component from HOMERUN is lower than that from the [S II] doublet and HOMERUN requires a contribution of more than 70 from the SF component to reproduce the [S II] emission (see Table 3). This highlights the strength of our decomposition method: while classical [S II] diagnostics measure a global electron density, HOMERUN disentangles the different ionizing sources and shows that the SF component alone has a much lower density. Finally, the derived ionization parameters are log(U) = -1.4 for the AGN component and a lower log(U) of -3.3 for the SF component, both in line with values found in H II regions (e.g. Kewley et al., 2001; Dopita et al., 2006) and AGN narrow-line regions (e.g. Groves et al., 2004). This supports the effectiveness of our decomposition in isolating distinct ionization regimes.
@@
Figure 7 shows the results of the fit. The left panel shows the behavior of the loss function as a function of oxygen abundance. The minimum of the curve is found for the minimum loss function of = 0.6, corresponding to 12+log(O/H) = 8.91. The left panels represent the weights of the single-cloud models with different U and NH which were derived for the best fit, for the AGN and SF components in the upper and lower panel, respectively. Using the weights, it is possible to compute the average density and ionization parameter of the clouds, which are identified by the black filled diamond. Only a small number of single-cloud models have non-zero weights, and the AGN and SF components contribute differently across the various lines, depending on the ionization state. In particular, as indicated in Tables 3 and 4, high-ionization lines (e.g., Ne V, O III) are predominantly reproduced by the AGN component, while lower-ionization lines (e.g., N II, Cl II, Ar II) are mainly explained by the SF component. Interestingly, the H luminosity is produced more than 70 by AGN excitation.
@@
Using the same procedure, we fitted with HOMERUN the emission lines in the spectra extracted from the regions marked in Fig. 1 (see Figs. 2 and 9). In the outflow-dominated region (green spectrum), we find for the SF component = 2.8, / cm-3) = 0.9, and = -3.4, and for the AGN component = 1.4, / cm-3) = 3.3, and = -1.4. In the stellar disk region (blue spectrum), we find for the SF component = 2.9, / cm-3) = 1.4, and = -3.3, and for the AGN component = 1.5, / cm-3) = 5.5, and = -1.5. These values are consistent with those obtained from the spectrum extracted within a 1.5″ radius aperture (see Fig. 6). Moreover, SF emission accounts for 64 of the H luminosity in the outflow region and 88 in the stellar disk region.
@@ 3.5.4 Ionized gas mass from HOMERUN
A key outcome of the HOMERUN fitting procedure is the ability to relate the luminosity of an emission line to the mass of the ionized gas associated. Since the fit is based on a physically consistent combination of AGN and SF single-cloud models, it allows us to disentangle the contributions from the different ionizing sources and connect the observed emission to intrinsic gas properties. In this way, under the assumption that the outflowing gas is fully ionized by the AGN, we avoided including the contribution from systemic or SF-related components, ensuring that the derived masses trace the outflowing gas itself. In particular, for each component, HOMERUN provides both the intrinsic (i.e., extinction-corrected) luminosities (see Tables 3 and 4) and the corresponding ionized gas mass. From the underlying single-cloud CLOUDY models, one obtains for each model a surface brightness in H (i.e. ) and a gas mass surface density (i.e. ). These model outputs define a direct proportionality between H luminosity and ionized gas mass. By rescaling this relation to the total luminosity of a specific emission line (e.g., [Ne V] or [O III]) and taking into account the AGN fraction as determined from the HOMERUN fit, we derive a general expression:
@@
where and are the intrinsic line luminosities emitted by the AGN component, and and are the total observed luminosities (i.e. without the HOMERUN separation between AGN and SF). The scaling factors applied to the observed luminosities account for both the fractional contribution of the AGN component and the effect of reddening. The scaling factors of in Equations 2 and 4 refer to extinction-corrected luminosities and can be directly compared with the standard conversion factors adopted in analytical calculations of the ionized gas masses (Equations 10 and 11, see also the discussion in the next section).
We compute an outflow mass of = (109 2) 103 and = (88 9) 103 within the 1.5″ radius aperture. These differences in mass, obtained with HOMERUN, reflect the discrepancy between the observed and model-predicted line ratios. Indeed, the observed ratio between [O III] and [Ne V] fluxes is 0.81 times the model ratio (see Appendix F), which matches the ratio between the [O III]- and [Ne V]-based outflow masses. We emphasize that the scaling relations in Equations 2-5 for the ionized outflow masses are not universally applicable to other AGN. They are derived specifically from the HOMERUN multi-cloud modeling customized for the physical conditions and emission properties of NGC 1365. Therefore, applying these relations to other galaxies requires performing a dedicated HOMERUN fit tailored to the individual characteristics of each galaxy.
@@ 3.6 Outflow energetics
In the previous section, with HOMERUN we obtain an outflow mass of = (109 2) 103 and = (88 9) 103 within a 1.5″ radius aperture centered on the outflow region (see Fig. 6 and Section 3.5.3). In addition, following Carniani et al. (2015), in Appendix G we present a detailed derivation of the outflow mass of [Ne V] and [O III]. Overall, with the classic method we find in the same spatial region = (2.6 1.0) 103 and = (20 7) 104 . The large difference between the HOMERUN and classical mass estimates (factors of 40 for [Ne V] and 2 for [O III]) clearly highlights the fundamental limitations of the standard approach. The classical method relies on several strong assumptions, such as considering that all the species are in the ionization state responsible for the observed line (e.g. O2+ for [O III], Ne4+ for [Ne V]), adopting a single average density (from [S II] in the optical and [Ne V] in the mid-IR), and correcting for attenuation using the Balmer decrement. In contrast, HOMERUN does not require assumptions about ionization levels, temperature, or metallicity, and accounts for differential reddening between the two components. The reliability of the more complex HOMERUN approach is demonstrated by the fact that classical method mass estimates differ by a factor of 75 between different tracers, highlighting the inherent inconsistencies of the traditional approach. Indeed, for example, the HOMERUN fitting shows that the fraction of neon in the form of Ne4+ in the best-fit models is only 10, while for O2+ the fraction is significantly higher, around 60 of the total oxygen. This is certainly one of the main factors contributing to the large discrepancy between the two mass estimates. Moreover, it should be noted that [O III] and [Ne V], having different critical densities and luminosities, trace gas from slightly different physical regions within the outflowing clouds, which further contributes to the observed differences in mass estimates beyond the methodological inconsistencies of the classical approach.
We assume that the flux density within each spaxel is constant over time and that the outflow subtends a solid angle . Therefore, we compute the mass outflow rate as , i.e. the amount of ionized mass crossing a distance with velocity . We also calculate the outflow kinetic energy as and the outflow power as .
@@
We infer the [Ne V] and [O III] intrinsic (i.e. de-projected) radial velocities of the outflow (vout) using the 3D multi-cloud kinematic model (Marconcini et al., 2023). This model has demonstrated exceptional power in reproducing outflow features with unprecedented detail, leveraging IFU data from a variety of instruments (Marconcini et al., 2023; Cresci et al., 2023; Perna et al., 2024; Ulivi et al., 2025; Marconcini et al., 2025a, b). To reproduce the outflow features in NGC 1365, we adopted a conical outflow morphology with maximum extension of 5 arcsec ( 470 pc) and position angle333The position angle is measured clockwise from North. of 235∘. The free parameters of the fit are the inclination of the cone axis with respect to the line of sight () and the intrinsic outflow velocity (vout). With we reproduce the line emission on a spaxel-by-spaxel basis, which as a consequence guarantee that the model reproduces the observed moment maps, with discrepancies between observed and modeled moment maps below 5. We found that the best-fit parameters that reproduce the [Ne V] outflow features in NGC 1365 are = 82 6∘, and vout = 390 35 km s-1, consistent with the results in Marconcini et al. (2025b). Fig. 8 shows the 3D representation of this conical outflow. Moreover, we used to fit the [O III] spatially resolved emission assuming the same outflow geometry as for the [Ne V] emission and found = 84 5∘, and vout = 470 40 km s-1. These results are also valid in the region where we extracted the spectrum used for the HOMERUN fitting.
So, adopting R = 284 pc ( 3″, i.e. the sky-projected extension of the 1.5″ radius aperture), we obtain the results of the outflow energetics reported in Table 1. From HOMERUN, we calculate consistent mass outflow rates for [Ne V] and [O III], namely = M⊙ yr-1 and = M⊙ yr-1, respectively. Instead, with the classic approach the [O III]-derived outflow rate is two orders of magnitude larger than that from [Ne V]. This discrepancy is primarily driven by the underestimation of the [Ne V] outflow mass in the classical approach. Specifically, the classical [Ne V] outflow mass is 75 times smaller than the [O III] estimate and 40 times smaller than the HOMERUN value, while the classical [O III] mass is only a factor of 2 higher than the HOMERUN mass. This asymmetric bias clearly demonstrates that the limitations of the standard method predominantly affects high-ionization tracers like [Ne V], leading to severely underestimated masses and consequently unreliable mass outflow rate calculations.

### v4

@@ 3.5 Photoionization modeling with HOMERUN
We modeled the full set of optical to mid-IR emission using state-of-the-art photoionization models. Specifically, we employed the Highly Optimized Multi-cloud Emission-line Ratios Using photo-ionizatioN (HOMERUN) modeling framework developed by Marconi et al. (2024).
3.5.1 Model grids of AGNs and SF
@@
Specifically, the shape of the AGN ionizing radiation field is described by a power law with UV slope , an exponential cutoff exp(-h/k TMax) and the X-ray component slope of =-1.0 linked to the UV through the . We have computed models with the following combinations of log(TMax/K) = 4.0, 4.5, 5.0, 5.5, 6.0, 6.5, 7.0 and .
As is described in Marconi et al. (2024), the incident spectra for the SF model grids are stellar population models from BPASS v2.3 (Byrne et al., 2022; Stanway and Eldridge, 2018) including binary stellar evolution. These models use a Kroupa (2001) initial mass function and an upper mass cutoff of 300 M⊙. We selected models with solar value for the stellar metallicity log(Zstar)=-1.7 and ages in log(age/Myr) of 6.0, 6.4, 6.6, 7.1, 8.8.
The other model parameters are identical for the two suites of models (AGN and SF); namely, the ionization parameter log(U) (i.e., the number of ionizing photons compared to that of atoms of neutral hydrogen) ranges from to in step of 0.5 and the hydrogen density of the gas, log(nH / cm-3) from 0 to 7 in steps of 1. For the first iteration HOMERUN spans metallicities log(Z/Z⊙) from -1.0 to 0.4 in steps of 0.2. After this first iteration, HOMERUN performs a final refinement in steps 0.02 on predictions interpolated across the finer grid.
3.5.2 HOMERUN modeling approach
@@
As is explained in detail in Marconi et al. (2024), the fitting process starts by selecting a grid of CLOUDY models, defined by hydrogen density () and ionization parameter () for a fixed ionizing spectrum (), gas phase oxygen abundance (), and elemental abundance ratios (). For each grid, the code searches for the best-fitting multicloud model by combining individual single-cloud templates. This is achieved by solving a non-negative least squares problem, whereby the weights of each cloud are constrained to be non-negative. The procedure is conceptually analogous to stellar population fitting methods (e.g., pPXF; Cappellari and Emsellem 2004), where a galaxy spectrum is reproduced as a linear combination of stellar templates with positive weights. No regularization is applied in the minimization, allowing for full flexibility in the weight distribution. The process is repeated across all available grids, spanning different ionizing continua, metallicities, and abundance scalings. The optimal solution is identified as the one minimizing the loss function , which quantifies the deviation between model predictions and observed emission-line fluxes. This loss function is effectively a reduced statistic, so values indicate a statistically good fit.
When exploring different metallicities, elemental abundances were scaled from the solar photospheric values of Asplund et al. (2021), except for carbon and nitrogen, which were rescaled following the prescriptions of Nicholls et al. (2017). During the fit, all elemental abundances except oxygen were allowed to vary from their initial values. Scaling factors were used for the AGN and SF components. Since the AGN component was assumed to be dust-free, while the SF component includes dust, refractory elements can have different scaling factors (due to depletion), whereas non-refractory elements share the same scaling factor in both components.
@@ 3.5.3 HOMERUN fitting results
We fit the fluxes of 60 ionized and neutral atomic emission lines measured from the MIRI and MUSE spectra (32 from MIRI and 28 from MUSE, see Figs. 2 and 9) extracted within a 1.5″ radius aperture centered on the outflow region (see Fig. 6). This aperture was chosen as the largest circular region that is fully covered by all MIRI channels, ensuring the detection of the maximum number of emission lines. The unique approach of HOMERUN enables us to reproduce at the same time 60 different emission lines in different ionization stages, with the majority of them reproduced with better than 15 accuracy. The data are best matched by a SF component with mean = 3.3, / cm-3) = 1.3, and = -3.3 and an AGN component with mean = 1.4, / cm-3) = 2.9, and = -1.4 (see Fig. 7). Being results of ionized gas region, in the following we will consider .
@@
We find that the SF component is more dust-attenuated than the outflow, in agreement with the extinction map obtained from the Balmer decrement (see Fig. 12) where the higher AV values trace the circumnuclear ring and the nucleus. Interestingly, the difference in AV between SF and AGN components is in line with the median values found in MAGNUM galaxies dividing the emission in disk and outflow (Mingozzi et al., 2019). In Section 3.4, we computed the electron density from the mid-IR [Ne V] and the optical [S II] line doublet ratios within the same 1.5″ radius aperture. The density derived from [Ne V], /cm-3) 2.87, is consistent with the average value of the AGN model components, /cm-3) = 2.9. This is expected since HOMERUN attributes the entire emission of [Ne V] to the AGN (see Table 4), as stars may not produce hard-enough radiation to account for the emission of this transition. Conversely, the [S II] lines yield a higher density of log(ne/cm-3) 2.5 within the 1.5″ radius aperture, compared to log(ne/cm-3) 1.3 from HOMERUN. This is because the average density of the SF component from HOMERUN is lower than that from the [S II] doublet and HOMERUN requires a contribution of more than 70 from the SF component to reproduce the [S II] emission (see Table 3). This highlights the strength of our decomposition method: while classical [S II] diagnostics measure a global electron density, HOMERUN disentangles the different ionizing sources and shows that the SF component alone has a much lower density. Finally, the derived ionization parameters are log(U) = -1.4 for the AGN component and a lower log(U) of -3.3 for the SF component, both in line with values found in H II regions (e.g., Kewley et al., 2001; Dopita et al., 2006) and AGN narrow-line regions (e.g., Groves et al., 2004). This supports the effectiveness of our decomposition in isolating distinct ionization regimes.
@@
Figure 7 shows the results of the fit. The left panel shows the behavior of the loss function as a function of oxygen abundance. The minimum of the curve is found for the minimum loss function of = 0.6, corresponding to 12+log(O/H) = 8.91. The left panels represent the weights of the single-cloud models with different U and NH which were derived for the best fit, for the AGN and SF components in the upper and lower panel, respectively. Using the weights, it is possible to compute the average density and ionization parameter of the clouds, which are identified by the green star. Only a small number of single-cloud models have non-zero weights, and the AGN and SF components contribute differently across the various lines, depending on the ionization state. In particular, as is indicated in Tables 3 and 4, high-ionization lines (e.g., Ne V, O III) are predominantly reproduced by the AGN component, while lower-ionization lines (e.g., N II, Cl II, Ar II) are mainly explained by the SF component. Interestingly, the H luminosity is produced more than 70 by AGN excitation.
@@
Using the same procedure, we fit with HOMERUN the emission lines in the spectra extracted from the regions marked in Fig. 1 (see Figs. 2 and 9). In the outflow-dominated region (green spectrum), we find for the SF component = 2.8, / cm-3) = 0.9, and = -3.4, and for the AGN component = 1.4, / cm-3) = 3.3, and = -1.4. In the stellar disk region (blue spectrum), we find for the SF component = 2.9, / cm-3) = 1.4, and = -3.3, and for the AGN component = 1.5, / cm-3) = 5.5, and = -1.5. These values are consistent with those obtained from the spectrum extracted within a 1.5″ radius aperture (see Fig. 6). Moreover, SF emission accounts for 64 of the H luminosity in the outflow region and 88 in the stellar disk region.
@@ 3.5.4 Ionized gas mass from HOMERUN
A key outcome of the HOMERUN fitting procedure is the ability to relate the luminosity of an emission line to the mass of the ionized gas associated. Since the fit is based on a physically consistent combination of AGN and SF single-cloud models, it allows us to disentangle the contributions from the different ionizing sources and connect the observed emission to intrinsic gas properties. In this way, under the assumption that the outflowing gas is fully ionized by the AGN, we avoided including the contribution from systemic or SF-related components, ensuring that the derived masses trace the outflowing gas itself. In particular, for each component, HOMERUN provides both the intrinsic (i.e., extinction-corrected) luminosities (see Tables 3 and 4) and the corresponding ionized gas mass. From the underlying single-cloud CLOUDY models, one obtains for each model a surface brightness in H (i.e., ) and a gas mass surface density (i.e., ). These model outputs define a direct proportionality between H luminosity and ionized gas mass. By rescaling this relation to the total luminosity of a specific emission line (e.g., [Ne V] or [O III]) and taking into account the AGN fraction, , as determined from the HOMERUN fit, we derive a general expression:
@@
where and are the intrinsic line luminosities emitted by the AGN component, and and are the total observed luminosities (i.e., without the HOMERUN separation between the AGN and SF). The scaling factors applied to the observed luminosities account for both the fractional contribution of the AGN component and the effect of reddening. The scaling factors of in Equations 2 and 4 refer to extinction-corrected luminosities and can be directly compared with the standard conversion factors adopted in analytical calculations of the ionized gas masses (Equations 10 and 11, see also the discussion in the next section).
We computed an outflow mass of = (109 2) 103 and = (88 9) 103 within the 1.5″ radius aperture. These differences in mass, obtained with HOMERUN, reflect the discrepancy between the observed and model-predicted line ratios. Indeed, the observed ratio between [O III] and [Ne V] fluxes is 0.81 times the model ratio (see Appendix F), which matches the ratio between the [O III]- and [Ne V]-based outflow masses. We emphasize that the scaling relations in Equations 2-5 for the ionized outflow masses are not universally applicable to other AGNs. They are derived specifically from the HOMERUN multicloud modeling customized for the physical conditions and emission properties of NGC 1365. Therefore, applying these relations to other galaxies requires performing a dedicated HOMERUN fit tailored to the individual characteristics of each galaxy. In the next section, we compare the outflow mass obtained with HOMERUN modeling with that estimated following a standard and commonly used approach (e.g., Cano-Díaz et al. 2012; Carniani et al. 2015, see Appendix G and Table 1).
@@ 3.6 Outflow energetics
In the previous section, with HOMERUN we obtain an outflow mass of = (109 2) 103 and = (88 9) 103 within a 1.5″ radius aperture centered on the outflow region (see Fig. 6 and Section 3.5.3). In addition, following Carniani et al. (2015), in Appendix G we present a detailed derivation of the outflow mass of [Ne V] and [O III]. Overall, with the classic method we find in the same spatial region = (2.6 1.0) 103 and = (20 7) 104 . The large difference between the HOMERUN and classical mass estimates (factors of 40 for [Ne V] and 2 for [O III]) clearly highlights the fundamental limitations of the standard approach. The classical method relies on several strong assumptions, such as considering that all the species are in the ionization state responsible for the observed line (e.g., O2+ for [O III], Ne4+ for [Ne V]), adopting a single average density (from [S II] in the optical and [Ne V] in the mid-IR), and correcting for attenuation using the Balmer decrement. In contrast, HOMERUN does not require assumptions about ionization levels, temperature, or metallicity, and accounts for differential reddening between the two components. The reliability of the more complex HOMERUN approach is demonstrated by the fact that classical method mass estimates differ by a factor of 75 between different tracers, highlighting the inherent inconsistencies of the traditional approach. Indeed, for example, the HOMERUN fitting shows that the fraction of neon in the form of Ne4+ in the best-fit models is only 10, while for O2+ the fraction is significantly higher, around 60 of the total oxygen. This is certainly one of the main factors contributing to the large discrepancy between the two mass estimates. Moreover, it should be noted that [O III] and [Ne V], having different critical densities and luminosities, trace gas from slightly different physical regions within the outflowing clouds, which further contributes to the observed differences in mass estimates beyond the methodological inconsistencies of the classical approach.
We assume that the flux density within each spaxel is constant over time and that the outflow subtends a solid angle, . Therefore, we computed the mass outflow rate as , i.e., the amount of ionized mass crossing a distance, , with velocity, . We also calculated the outflow kinetic energy as and the outflow power as .
@@
We inferred the [Ne V] and [O III] intrinsic (i.e., de-projected) radial velocities of the outflow (vout) using the 3D multicloud kinematic model (Marconcini et al., 2023). This model has demonstrated exceptional power in reproducing outflow features with unprecedented detail, leveraging IFU data from a variety of instruments (Marconcini et al., 2023; Cresci et al., 2023; Perna et al., 2024; Ulivi et al., 2025; Marconcini et al., 2025a, b). To reproduce the outflow features in NGC 1365, we adopted a conical outflow morphology with maximum extension of 5 arcsec ( 470 pc) and position angle333The position angle is measured clockwise from north. of 235∘. The free parameters of the fit are the inclination of the cone axis with respect to the line of sight () and the intrinsic outflow velocity (vout). With we reproduce the line emission on a spaxel-by-spaxel basis, which as a consequence guarantee that the model reproduces the observed moment maps, with discrepancies between observed and modeled moment maps below 5. We found that the best-fit parameters that reproduce the [Ne V] outflow features in NGC 1365 are = 82 6∘, and vout = 390 35 km s-1, consistent with the results in Marconcini et al. (2025b). Fig. 8 shows the 3D representation of this conical outflow. Moreover, we used to fit the [O III] spatially resolved emission assuming the same outflow geometry as for the [Ne V] emission and found = 84 5∘, and vout = 470 40 km s-1. These results are also valid in the region where we extracted the spectrum used for the HOMERUN fitting.
So, adopting R = 284 pc ( 3″, i.e., the sky-projected extension of the 1.5″ radius aperture), we obtained the results of the outflow energetics reported in Table 1. From HOMERUN, we calculated consistent mass outflow rates for [Ne V] and [O III], namely = M⊙ yr-1 and = M⊙ yr-1, respectively. Instead, with the classic approach the [O III]-derived outflow rate is two orders of magnitude larger than that from [Ne V]. This discrepancy is primarily driven by the underestimation of the [Ne V] outflow mass in the classical approach. Specifically, the classical [Ne V] outflow mass is 75 times smaller than the [O III] estimate and 40 times smaller than the HOMERUN value, while the classical [O III] mass is only a factor of 2 higher than the HOMERUN mass. This asymmetric bias clearly demonstrates that the limitations of the standard method predominantly affects high-ionization tracers like [Ne V], leading to severely underestimated masses and consequently unreliable mass outflow rate calculations.